Feb. 25, 2011

Members: Leigh, Mark

Mark arrive: 7:30 pm, Leigh arrive: 8:30

Temp: -8 C

Weather clear with some moisture at hight elevation.

Mark & Leigh set up Mark’s Tak on EQ6 and Leigh set up Vixen 95mm on HEQ5 on outdoor pad. Mark took images of M101 and M42 with DSLR. Leigh tried out new laptop to test battery durability and whether programs set up properly. Leigh was able to control telescope mount with laptop, however the cameras would not be found. Probably do not have correct drivers yet for Windows7 64bit system. Will need to do more work in warmth of own home. Laptop battery performed as advertised. Ran 4.5 hours and still showed 4 hours remaining. Happy!

Liquid mirror telescope working tonight. Laser visible and showed incoming clouds by 12:15am. Bands of cloud moved throug between 1:00 and 1:30am.

Departed: 1:30 am

Temp: -7C

Darkness meter: n/a